OpenStudy (anonymous):

A partial solution set is given for the equation. Find the complete solution set. 2x3 − 5x2 − 14x + 8 = 0; {1/2}

myininaya (myininaya):

do you know synetheic division? 1/2| 2 -5 -14 8 | 1 -2 -8 2 -4 -16 |0 2x^2-4x-16=0 try to solve this for x

myininaya (myininaya):

divide both sides by 2 x^2-2x-8=0 (x-4)(x+2)=0 x=4 x=-2